# MathGAP: Out-of-Distribution Evaluation on Problems with Arbitrarily Complex Proofs

### v1

@@ Abstract
Large language models (LLMs) can solve arithmetic word problems with high accuracy, but little is known about how well they generalize to problems that are more complex than the ones on which they have been trained. Empirical investigations of such questions are impeded by two major flaws of current evaluations: (i) much of the evaluation data is contaminated, in the sense that it has already been seen during training, and (ii) benchmark datasets do not capture how problem proofs may be arbitrarily complex in various ways. As a step towards addressing these issues, we present a framework for evaluating LLMs on problems that have arbitrarily complex arithmetic proofs, called MathGAP. MathGAP generates problems that follow fixed proof specifications—along with chain-of-thought reasoning annotations—enabling systematic studies on generalization with respect to arithmetic proof complexity. We apply MathGAP to analyze how in-context learning interacts with generalization to problems that have more complex proofs. We find that among the models tested, most show a significant decrease in performance as proofs get deeper and wider. This effect is more pronounced in complex, nonlinear proof structures, which are challenging even for GPT-4o. Surprisingly, providing in-context examples from the same distribution as the test set is not always beneficial for performance. In particular, zero-shot prompting as well as demonstrating a diverse range of examples that are less complex than the test data sometimes yield similar or higher accuracies.
@@ 1 Introduction
High performance on reasoning benchmarks are often taken as evidence that transformer-based large language models (LLMs) are able to “reason”. However, many current evaluations are unreliable since it is likely that the problems they contain are present in the model’s training data (Sainz et al., 2023; Deng et al., 2024; Zhang et al., 2024). Moreover, most evaluations fail to capture that reasoning problems can be arbitrarily complex, through composition of subproofs and use of multiple rules of inference. High accuracy on a specific set of problems in a benchmark dataset is therefore not sufficient to conclude that LLMs can generalize to more complex, unseen problems. To obtain a more appropriate empirical measure of reasoning ability, one must evaluate on data that is not present in any benchmark dataset, containing proofs that are more complex than those used in training and context window alike.
Math word problems (MWPs) are one of the most frequently used testbeds for LLM reasoning. Yet there exist few, if any, systematic analyses on out-of-distribution (OOD) generalization in regards to proof complexity that avoid contamination in this domain. We suspect that this is in part due to a lack of formalism for characterizing proof complexity, which we argue is necessary for making precise claims about generalization behavior for reasoning problems.
@@
In this paper, we present a framework for evaluating Mathematical Generalization on Arithmetic Proofs—MathGAP. First, we discuss how to characterize solutions to MWPs in terms of proof trees, in which the nodes are labeled with logical statements about the problem under a world-model framework (Opedal et al., 2023) and the edges are induced by proof steps on such logical statements. Doing so reveals several ways of characterizing a problem’s complexity based on its proof tree, e.g., (non)linearity, depth, width, and the ordering of nodes (all defined in § 3.2). With this formalism, we develop a method to generate problems with specified proof tree characteristics. Because this method is based on proof trees, it also enables rich annotations in the form of ground-truth reasoning traces. More specifically, traversing the proof trees in post order, we obtain chain-of-thought (CoT; Wei et al., 2022) reasoning traces in an automated manner. Fig. 1 illustrates the generation method, showing a nonlinearauthor=bernhard,color=white!40,size=,fancyline,caption=,]does everybody know what nonlinear means in this context, or should this be defined? Could mention that this will be defined below, at the end of Sec 3  andreas: good point, added above. I am now worried that (non)linearity is not a standard term here. Maybe we should rebrand it to something like “proof tree topology”? Would be more general as well proof tree with depth $2$ and width $5$, in which the ordering of the sentences is given by traversing the leaf nodes in left-to-right order. This generation method forms the backbone of MathGAP, our framework for studying OOD generalization gaps on arithmetic proofs. MathGAP avoids data contamination since the distribution of generated test problems can always be different from those in training or in context—by design. When the performance on problems at one level of complexity hits saturation, we can flexibly generate a new set of problems which are even more complex. This is in spirit similar to Dynabench (Kiela et al., 2021) but, importantly, does not require human annotators.
@@
MathGAP opens up the possibility for several new empirical investigations. [^1] Code to generate problems with MathGAP will be publicly released with the final version. In the present study, we perform a systematic analysis on whether LLMs can use simple examples in context to solve more complex ones at inference. To this end, we consider two different distributions of in-context CoT demonstrations: examples with only the simplest possible proof trees, and a range of examples with varying complexity, akin to curriculum learning (Bengio et al., 2009). We then evaluate on test sets containing problems that are increasingly more complex than the in-context examples. The performance is compared to two baselines of in-context demonstrations, a prompt that has in-distribution examples with respect to the test set, and a zero-shot prompt. Our study includes Mixtral-8x7B (Jiang et al., 2024), Llama3-8B, Llama3-70B (Llama Team, 2024), GPT-3.5-Turbo, GPT-4o (OpenAI, 2024), and a few additional evaluations on o1-preview. We present several findings:
  • As expected, there is consistent decrease in performance as proof depth and width increase. The decrease is particularly notable for nonlinear problems, for which we observe a general inability of the evaluated LLMs to solve complex problems, even for GPT-4o.
@@
  • Generalization across orderings appears to be different, however. We conduct experiments where we move sentences to the front of the problem, and find that the problem is easiest for LLMs if the sentence is moved from near the beginning or end, rather than from the middle of the problem. This is surprising since one would expect a monotonically decreasing relationship between accuracy and the distance of the movement.
@@
  • We see no clear relationship between in-context distribution and performance. Contrary to our expectations, in-distribution examples are not always beneficial for performance, as compared to zero shot. Demonstrating a range of examples with varying complexityauthor=bernhard,color=white!40,size=,fancyline,caption=,]sounds like in-context curriculum learning?  andreas: yes true, I’ve added some references to that. The difference is that we randomly sample the ordering of the examples; this difference is explained in § 5 is usually preferable to demonstrating simple examples, suggesting that LLMs benefit from a diverse prompt.
@@
  • The o1-preview model outperforms the other models on nonlinear problems. Yet, we are able to create a test set of which it solves only 5%—demonstrating that MathGAP is future-proof.
@@ 2 Related Work
Machine learning, broadly, is the study of algorithms that can learn from data, and generalize to unseen data. Machine learning models must therefore be evaluated on unseen test sets. As it turns out, modern-day LLMs are not being properly evaluated in this regard: Much of the data they are evaluated on has already been used during model training, both labeled (Dodge et al., 2021; Deng et al., 2024) and unlabeled (Elazar et al., 2023). Researchers have raised serious concerns about these issues (Jacovi et al., 2023; Sainz et al., 2023). For arithmetic reasoning specifically, Zhang et al. (2024) recently found evidence that the widely used GSM8k benchmark (Cobbe et al., 2021) is contaminated, and presented evaluations on a new dataset which is not publicly released. MathGAP mitigates data contamination by creating new, more complex synthetic test sets. author=mrinmaya,color=pink!40,size=,fancyline,caption=,]I am less sure about this. Maybe we should just say that we can use MATHGap to always create new test problems  andreas: Yeah I guess that’s a stronger argument, edited. However I think we should refrain from publishing our test data though. Instead only publish the data-generating code. Do you disagree?  abu: oh interesting, so keep the random seed for our test data hidden? i suppose people can still reproduce our results to an extent, with sampling noise  andreas: yes exactly, just so that models don’t overfit to our particular test examples
@@
Generalization to data from a different distribution (i.e., OOD data) is a core problem in machine learning, and has been studied extensively also in the context of LLM reasoning (Schwarzschild et al., 2021; Razeghi et al., 2022; An et al., 2023; Kudo et al., 2023; Liu et al., 2023; Saparov & He, 2023; Zhang et al., 2023; Thomm et al., 2024). While LLMs’ performance on OOD data can be improved by various techniques (Anil et al., 2022; Hu et al., 2024), generalization to problems of arbitrary complexity under a set of known inference rules is still an open problem. For instance, Dziri et al. (2023) demonstrate that the accuracy of transformer-based LLMs on compositional generalization rapidly approaches zero as the complexity of the problem increases—regardless of in-context guidance. Saparov et al. (2023) focus on theorem proving for logical reasoning, presenting a systematic analysis on generalization ability of LLMs with respect to proof size and diversity of inference rules. Our paper contributes to this literature by providing an evaluation framework to systematically study OOD generalization in regards to arithmetic proof complexity.
@@
It is common to treat MWPs as a testbed for studying reasoning abilities of LLMs (Cobbe et al., 2021; Patel et al., 2021; Fu et al., 2023; Shakarian et al., 2023; Stolfo et al., 2023; Zong & Krishnamachari, 2023); such problems are useful because solving them requires several distinct skills, yet, they remain conceptually simple (Stern, 1993). However, we are not aware of many studies that investigate generalization in regards to problem complexity in this domain. Our study presented in § 5 relates to Hase et al. (2024), who fine-tune LLMs on easy problems and evaluate them on harder ones at inference. Their complexity metrics focus on problem length, but do not capture how a longer reasoning problem gets more complex (i.e., the shape of the proof tree). Moreover, they evaluate on GSM8k, which is contaminated. Mishra et al. (2024) proposed a method that uses GPT-4 to generate new synthetic problems from symbolic specifications. Their approach is scalable, but using an LLM to generate problems may (i) introduce bias (Boguraev et al., 2024; Opedal et al., 2024) and (ii) produce problems that are unfaithful to their specifications. Mirzadeh et al. (2024) recently performed evaluations on a version of GSM8k in which the variable names and numbers have been altered, and found that LLMs are surprisingly sensitive to such substitutions. MathGAP is more general; while it allows for such substitutions, our main goal is to vary the proof structures.
Table 1: Examples of logical forms. A logical form consists of a predicate and a set of properties which is specific to that predicate, given here in separate columns. Each logical form induces a set of natural language sentences which preserve its semantics.
@@
partwhole (whole_agent=$\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}$, [$\texttt{a}_{1}$], …, and [$\texttt{a}_{n}$] combine the [f]s that they have. Alice and Bob combine the fruits that they have.
part_agent1=a1,
part_agentn=an,
whole_entity=f,
part_entity=e)
@@
3 A Formal Treatment of Math Word Problems
We aim to generate new problems of arbitrary complexity, so we require a formalism for characterizing complexity in a precise manner. First, we explain how the semantics expressed in math word problems can be written as sequences of logical forms (§ 3.1). We then show how to apply inference rules on these logical forms to deduce new ones (§ 3.2). This leads to a view of problem solutions as proof trees, the structure of which can be used to characterize the reasoning required to solve the problem.
@@ 3.1 Math word problems as logical forms
Each math word problem is represented as a sequence of logical forms under the formalism from Opedal et al. (2023; 2024). A logical form is a statement about a truth in the world that is being described by the problem, representing some arithmetic relationship between the number of entities possessed by one or several agents. Such statements can refer to the number of entities an agent has, how many more entities an agent has relative to another, the action of an agent giving their entities to another agent, etc. Formally, a logical form represents the semantics of a single sentence, and consists of a predicate that takes a set of properties as arguments. Every logical form has an agent property, an entity property and a quantity property. Other, optional, properties include unit and attribute. We define a world model as a sequence of logical forms that align with the sentences of a problem text, thus, representing the semantics of the entire problem. See Fig. 5 for the world model representation of the problem from Fig. 1.
@@
For a given predicate and set of properties, we write predicate(property1=x, property2=y, ...) to express a logical form. Property names are, however, often omitted for brevity, as in predicate(x, y, ...). There are two types of predicates: (i) those that represent ownership over some entity by an agent, called container (abbreviated cont), and (ii) those that represent arithmetic relationships between two sets of properties. The relationship predicates correspond to arithmetic concepts and are based on a taxonomy from the learning sciences (Riley et al., 1983); we use transfer , comparison (abbreviated comp), partwhole and comp-eq. [^2] The transfer and part-whole concepts are usually called change and combine, respectively, in learning science literature (Nesher et al., 1982). The comp-eq predicate was introduced in this paper in order to construct nonlinear problems of arbitrary length. Table 1 gives concrete example sentences for logical forms with each of these predicates. In words, a cont logical form expresses the fact that an agent owns a particular entity in some quantity (e.g., “Alice has 5 apples”). A comp logical form expresses how much more some agent owns of a particular entity than some other agent (e.g., “Bob has 3 fewer apples than Alice”). A transfer logical form expresses the transfer of a certain quantity of an entity from one agent to another (e.g., “Alice gave Bob 3 apples”). A partwhole logical form expresses a superset relation between the combined entities of several agents (e.g., Alice and Bob combine their fruits). Finally, a comp-eq logical form expresses that the quantity of a particular comp between two agents is equal to that of a comp between two other agents (e.g., “The difference between the number of apples that Alice and Bob have is the same as the difference between the number of apples that Charlie and David have”).
@@
The logical forms in the world model of a math problem are separated into the body and question. The question is a single logical form which appears at the end of the world model. It is characterized by having a placeholder variable instead of an explicit quantity, which corresponds to the correct answer of the problem. For example, the interrogative sentence “How many apples does Alice have?” is represented by the question cont(Alice, q, apples), where q is a placeholder variable. The body consists of all logical forms in the world model which are not questions. The answer to a problem is the logical form declared by the question with the correct numerical quantity substituted into its placeholder variable, although we will sometimes refer to only the correct numerical quantity as the answer as well.
@@
Table 2: Inference rule templates (left) with corresponding examples of rule applications in natural language (right). The variables a, b, c, d refer to agents, q, q1, q2 refer to quantities, and e, f refer to entities. Attribute and unit properties are excluded here, but analogous rules exist in which they are present. All inference rules are commutative except the transfer rule; see footnote 3. Rule applications of an inference rule with premises $L_{1},L_{2},\dots,L_{N}$ and consequent $L$ are written as $L^{\prime}_{1}.L^{\prime}_{2}.\,\dots\,.L^{\prime}_{N}.\vdash L^{\prime}.$, where $L^{\prime}$ is a natural language expression for the logical form $L$. Axioms are formed from the facts stated in the problem, as illustrated in Fig. 1.

### v3

@@ Abstract
Large language models (LLMs) can solve arithmetic word problems with high accuracy, but little is known about how well they generalize to more complex problems. This is difficult to study, as (i) much of the available evaluation data has already been seen by the most capable models during training, and (ii) existing benchmarks do not capture how problem proofs may be arbitrarily complex in various ways. In this paper, we present a data-generation framework for evaluating LLMs on problems with arbitrarily complex arithmetic proofs, called MathGAP. MathGAP generates problem statements and chain-of-thought reasoning traces according to specifications about their arithmetic proof structure, enabling systematic studies on easy-to-hard generalization with respect to complexity of proof trees. Using MathGAP, we find that LLMs show a significant decrease in performance as proofs get deeper and wider. This effect is more pronounced in complex, nonlinear proof structures, which are challenging even for the most capable models. The models are also sensitive to simple changes in sentence ordering. However, they remain capable of solving some complex problems, suggesting that reasoning generalization is noisy.
[Uncaptioned image]   https://github.com/eth-lre/mathgap-experiments
@@ 1 Introduction
High performance on reasoning benchmarks is often taken as evidence that transformer-based large language models (LLMs) can “reason”. However, many current benchmarks are unreliable, as it is likely that the problems they contain are present in the model training data (Sainz et al., 2023; Deng et al., 2024; Zhang et al., 2024). Moreover, most evaluations fail to capture that reasoning problems can be arbitrarily complex through composition of subproofs and use of multiple rules of inference. High accuracy on a specific set of problems in a benchmark dataset is therefore not sufficient to conclude that LLMs can generalize to more complex, unseen problems. To obtain a more appropriate empirical measure of reasoning ability, one must evaluate on data that is not present in any benchmark dataset, containing proofs that are more complex than the ones the model has already seen.
@@
In this paper, we aim to satisfy these desiderata. We consider math word problems (MWPs), which are one of the most frequently used testbeds for LLM reasoning, and propose a framework for evaluating Mathematical Generalization on Arithmetic Proofs—MathGAP. We start by discussing how solutions to MWPs can be characterized in terms of proof trees, in which nodes are labeled with logical statements about the problem under a world-model framework (Opedal et al., 2023), and the edges are induced by proof steps on such logical statements. Doing so reveals several ways of characterizing a problem’s complexity based on its proof tree, e.g., depth, width, shape, and the ordering of nodes (all defined in § 3.2). With this formalism, we develop a method to generate problems with specified proof tree characteristics. Because this method is based on proof trees, it also enables rich annotations in the form of ground-truth reasoning traces. More specifically, traversing the proof trees in post order, we obtain chain-of-thought (CoT; Wei et al., 2022) reasoning traces in an automated manner. Fig. 1 illustrates the generation method, showing a proof tree having depth $2$, width $5$, what we call nonlinear shape, and where the ordering of the sentences is given by traversing the leaf nodes in left-to-right order.
This generation method forms the backbone of MathGAP, our framework for studying out-of-distribution (OOD) generalization gaps on arithmetic proofs. MathGAP avoids data contamination since the distribution of generated test problems can always be different from those in training or in context, by design. When the performance on problems at one level of complexity hits saturation, we can flexibly generate a new set of problems that are even more complex. This is in spirit similar to Dynabench (Kiela et al., 2021) but, importantly, does not require human annotators.
@@
MathGAP enables many new empirical investigations. In this paper, we perform a systematic analysis on whether LLMs can use simple examples in context to solve more complex ones at inference. To this end, we consider two different distributions of in-context CoT demonstrations: examples with only the simplest possible proof trees, and a range of examples with varying complexity, akin to curriculum learning (Bengio et al., 2009). We then evaluate on test sets containing problems that are increasingly more complex than the in-context examples. The performance is further compared to in-distribution and zero-shot baselines. Our study includes Mixtral-8x7B (Jiang et al., 2024a), Llama3-8B, Llama3-70B (Llama Team, 2024), GPT-3.5-Turbo, GPT-4o (OpenAI, 2024), and a few additional evaluations on o1-preview and DeepSeek-R1 (DeepSeek-AI, 2025). We present several findings:
  1. i)
    Generalization to proof width is harder than to proof depth, but there is a steady decrease in performance as proofs get both deeper and wider. The performance decrease is particularly notable for the so-called nonlinear problems; solving them requires keeping intermediate steps longer in memory than for linear problems, for which intermediate steps can be consumed immediately (under post-order traversal). Even GPT-4o fails to solve the most complex nonlinear problems.
  2. ii)
    However, the performance decrease is moderate in most cases. This suggests that the models are able to generalize to some extent, albeit with a noisy form of reasoning (Prabhakar et al., 2024).
  3. iii)
@@
    All models are sensitive to a simple change in sentence ordering, in which one sentence is moved to the front of the problem. Moreover, we find that the problem is easiest for LLMs if the sentence is moved from the beginning or the end rather than from the middle of the problem. This is surprising since one would expect a monotonically decreasing relationship between accuracy and the distance of movement.
  4. iv)
@@
    We do not see a consistent advantage to using in-context examples. Contrary to our expectations, in-distribution examples are not always beneficial for performance, as compared to zero-shot. This suggests that some problems are hard to learn even when the LLM is given demonstrations of similar problems, independently of its generalization ability. Demonstrating a range of examples with varying complexity is usually preferable to demonstrating simple examples, suggesting that LLMs benefit from a diverse prompt.
  5. v)
@@
    o1 and R1 outperform the others on nonlinear problems. However, we are able to create a test set on which they solve only 5% and 11% respectively—indicating that MathGAP is future-proof.
@@ 2 Related Work
Machine learning, broadly, is the study of algorithms that can learn from data and generalize to unseen data. Machine learning models must therefore be evaluated on unseen test sets. As it turns out, modern-day LLMs are not being properly evaluated in this regard. Much of the data they are evaluated on has already been used during model training, both labeled (Dodge et al., 2021; Deng et al., 2024) and unlabeled (Elazar et al., 2024). Researchers have raised serious concerns about such data contamination issues (Jacovi et al., 2023; Sainz et al., 2023). For arithmetic reasoning specifically, Zhang et al. (2024) recently found evidence that the widely used GSM8k benchmark (Cobbe et al., 2021) is contaminated and presented evaluations on a new dataset that is not publicly released. MathGAP mitigates data contamination by creating new, more complex synthetic test sets.
@@
Generalization to data from a different distribution (i.e., OOD data) is a core problem in machine learning and has also been studied extensively in the context of LLM reasoning (Schwarzschild et al., 2021; An et al., 2023; Kudo et al., 2023; Liu et al., 2023; Saparov & He, 2023; Zhang et al., 2023; Mészáros et al., 2024; Thomm et al., 2024). While LLMs’ performance on OOD data can be improved by various techniques (Anil et al., 2022; Borazjanizadeh & Piantadosi, 2024; Hu et al., 2024), generalization to problems of arbitrary complexity under a set of known inference rules is still an open problem. For instance, transformer-based LLMs struggle to generalize to longer problem proofs when solving logical reasoning tasks (Dziri et al., 2023; Saparov et al., 2023). We contribute to this literature by providing an evaluation framework for OOD generalization to complex arithmetic proofs. Our work relates to scalable oversight (Bowman et al., 2022) and easy-to-hard generalization (Burns et al., 2023; Hase et al., 2024; Sun et al., 2024), which can be studied in a principled manner with MathGAP by characterizing what is easy and what is hard.
@@
MWPs are often treated as a testbed for studying reasoning abilities of LLMs (Patel et al., 2021; Fu et al., 2023; Shakarian et al., 2023; Stolfo et al., 2023; Zong & Krishnamachari, 2023; Huang et al., 2024; Ye et al., 2024; Yu et al., 2024); such problems are interesting because they require several distinct skills to solve, while remaining conceptually simple (Stern, 1993). However, we are not aware of much work on generalization in regard to problem complexity in this domain. Our study in § 5 is related to Hase et al. (2024), who fine-tune LLMs on easy problems and evaluate them on harder ones at inference, using GSM8k and other possibly contaminated datasets. Their complexity metrics focus on problem length, but do not capture how a reasoning problem gets more complex (i.e., through the characteristics of the proof tree). Mirzadeh et al. (2025) recently performed evaluations on a version of GSM8k in which variable names and numbers have been altered, and found that LLMs are surprisingly sensitive to such substitutions. MathGAP is more general; while it allows for such substitutions as well, our main goal is to vary the proof structures.
3 A Formal Treatment of Math Word Problems
We aim to generate new problems of arbitrary complexity, so we require a formalism to characterize complexity in a precise manner. First, we explain how the semantics expressed in MWPs can be written as sequences of logical forms (§ 3.1). We then show how to apply the inference rules on these logical forms to deduce new ones (§ 3.2). This leads to a view of problem solutions as proof trees, the structure of which can be used to characterize the reasoning required to solve the problem.
Table 1: Example logical forms. A logical form comprises a predicate and a set of properties specific to that predicate. Each logical form induces natural language sentences that preserve its semantics.
@@ 3 A Formal Treatment of Math Word Problems
partwhole (whole_agent=$\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}$, [$\texttt{a}_{1}$], …, and [$\texttt{a}_{n}$] have [q] [e]s combined.
quantity=q, Alice and Bob have 8 fruits combined.
whole_entity=e)
@@ 3.1 Math word problems as logical forms
Each MWP is represented as a sequence of logical forms under the formalism from Opedal et al. (2023; 2024), which is a subset of first-order logic. A logical form is a truth statement about the world that is being described by the problem, representing some arithmetic relationship between the number of entities possessed by one or several agents. Such statements can refer to the number of entities an agent has, how many more entities an agent has relative to another, the action of an agent giving their entities to another agent, etc. Formally, a logical form represents the semantics of a single sentence and consists of a predicate that takes a set of properties as arguments. Every logical form has an agent property, an entity property, and a quantity property. Other, optional properties include unit and attribute. We define a world model as a sequence of logical forms that align with the sentences of a problem text, thus representing the semantics of the whole problem.
@@
For a given predicate and set of properties, we write predicate(property1=x, property2=y, ...) to express a logical form. Property names are, however, often omitted for brevity, as in predicate(x, y, ...). There are two types of predicates: (i) those that represent the ownership of some entity by an agent, called container (abbreviated cont), and (ii) those that represent arithmetic relationships between two sets of properties. The relationship predicates correspond to arithmetic concepts and are based on a taxonomy from the learning sciences (Riley et al., 1983); we use transfer , comparison (abbreviated comp), partwhole and comp-eq. [^1] The transfer and part-whole concepts are usually called change and combine, respectively, in learning science literature (Nesher et al., 1982). The comp-eq predicate was introduced in this paper in order to construct nonlinear problems of arbitrary length—see App. D for details. Table 1 gives concrete example sentences for logical forms with each of these predicates, Fig. 5 shows the world model representation of the problem in Fig. 1, and App. A provides further intuition. [^2] In order to denote that the items of several agents are combined under a partwhole, we treat the conjunction of agents $\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}=\texttt{a}$ as another agent in Table 1. The same is done for entities, as shown in § 3.1.
@@
The logical forms in the world model of a problem are separated into body and question. The question is a single logical form that appears at the end of the world model. It holds a placeholder variable instead of an explicit quantity, which represents the correct answer to the problem. For example, the interrogative sentence “How many apples does Alice have?” is represented by the question cont(Alice, q, apples), where q is a placeholder variable. The body consists of all logical forms in the world model that are not questions. The answer to a problem is the logical form declared by the question with the correct numerical quantity substituted into its placeholder variable.
@@
Table 2: Inference rule templates (left) with corresponding examples of rule applications in natural language (right). The variables a, b, c, d refer to agents, q, q1, q2 refer to quantities, and e, f refer to entities. Attribute and unit properties are excluded here, but analogous rules exist in which they are present. All inference rules are commutative except the transfer rule; see footnote 4. Rule applications of an inference rule with premises $L_{1},L_{2},\dots,L_{N}$ and consequent $L$ are written as $L^{\prime}_{1}.L^{\prime}_{2}.\,\dots\,.L^{\prime}_{N}.\vdash L^{\prime}.$, where $L^{\prime}$ is a natural language expression for the logical form $L$. Axioms are formed from the facts stated in the problem, as illustrated in Fig. 1.